1/5 is greater than 3/16, but why? How do I do the math to figure out the answer myself? The problem that was given was: which number is smallest? A. 1/5 or B. 3/16 ? and I saw the answer but I want to know how they came up with that.
Answer by scott8148(6628)

make the denominators of the fractions the same , and then compare the numerators

the LCD is 80

1/5 = 16/80 ___ multiplying numerator and denominator by 16

3/16 = 15/80 ___ multiplying numerator and denominator by 5